OpenStudy (anonymous):

What is the equation of the line (in slope-intercept form) that passes through the given point and is perpendicular to the graph of the given equation? (6,negative -2); y = negative-3x + 4 y = negative-1 over 3x − 4 y = 1 over 3x + 4 y = negative-3x − 4 y = 1/3x − 4

OpenStudy (anonymous):

@gorv @Joel_the_boss

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

y = -3x + 4 can you tell me the slope of this line ?

OpenStudy (anonymous):

-3x

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position. So what is the slope of the given equation

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

correct....-3 is the slope. However, we are looking for a perpendicular line and that contains a negative reciprocal slope. So we need the negative reciprocal of -3. To find the negative reciprocal, just " flip " the slope and change the sign. So our new slope is 1/3....see how I flipped the slope and changed the sign. Now we use y = mx + b form slope(m) = 1/3 (6,-2)...x = 6 and y = -2 we already know the slope(m) so we have to find the y intercept(b) now we sub -2 = 1/3(6) + b -2 = 2 + b -2 - 2 = b -4 = b so your perpendicular line is : y = 1/3x - 4 any questions ?

OpenStudy (anonymous):

Thanks, now I understand :)

OpenStudy (texaschic101):

glad to help :)